An ideal step-down transformer has a primary coil of 400 turns and a secondary coil of 25 turns. Its primary coil is plugged into an outlet with 110 V(AC), from which it draws an rms current of 0.83 A.

1.)What is the voltage in the secondary coil?

2.)Calculate the rms current in the secondary coil.

3.)Assuming that the transformer secondary is driving a resistive load, calculate the average power dissipated in the resistor.
Answer #1

Np =400 , Ns = 25, Vp(rms) = 110 V , iP (rms) =0.83 A

(1) Is/Ip = Np/Ns

Is/ 0.83 = 400/25

Is = 13.28 A

(2) Pav = Ip*Vp = 110*0.83

Pav = 91.3 W